# Interpretable representation learning of quantum data enabled by probabilistic variational autoencoders

### v2

@@ III.1.2 LR-TFIM
As a second benchmark, we consider the long-range transverse field Ising model (LR-TFIM), which has been recently performed experimentally via trapped-ion simulators [31, 32, 17]. This model is a modification of the conventional TFIM in which spin–spin interactions decay algebraically with distance. Its Hamiltonian is given by
@@
In contrast, the latent space learned by the cpVAE (Fig. 3c) exhibits a markedly richer structure. The first latent neuron, $\mu_{1}$, correlates strongly with a magnetization-like observable (see Appendix D for more details), which is consistent with it capturing the dominant factor of variation in the data. The second, $\mu_{2}$, encodes a distinct feature that highlights the presence of an intermediate phase emerging at low values of $\alpha$ and $h$. To investigate this further, we applied symbolic regression [33] to the latent values of $\mu_{2}$, seeking a compact analytical expression $f(x)$ that best describes its behavior given the corresponding spin configuration $x$. This approach suggests that the neuron was encoding a quantity similar to the structure factor [34]
@@
As shown in Fig. 3(d), the structure factor, evaluated here at fixed momentum $k=0.8$ and spatial reference site $i=0$, highlights the same region of the phase diagram as captured by the latent variable $\mu_{2}$. From a physical perspective, the structure factor reveals a distinct intermediate regime between the ferromagnetic phase and the high $\beta$ paramagnetic regime. This feature, visible only at the chain boundaries (i.e., $i=0$ and $i=N$), suggests the emergence of a boundary-induced phase, likely arising from finite-size and open-boundary effects [35]. That the cpVAE identifies this behaviour without supervision underscores its potential as a tool for uncovering nontrivial structure in many-body systems and guiding the formulation of new hypotheses.
@@ III.2 Experimental Rydberg Atoms
As a final benchmark, we train the cpVAE with experimental snapshots obtained from a programmable Rydberg atom array [16]. The setup consists of a two-dimensional square lattice with $N=L\times L$ atoms, where each atom resides in either its electronic ground state |$\ket{g}$ or an excited Rydberg state |$\ket{r}$. The latter is achieved via laser-driven excitation of an outer electron to a high-lying energy level, giving rise to strong, long-range interactions between excited atoms. This platform provides an ideal testbed for the cpVAE, as it exhibits a variety of quantum phenomena, including frustration, quantum criticality and emergent spatial order, all encoded in the projective measurement snapshots. Our aim here is to assess whether cpVAE can extract meaningful latent representations from such data, despite the presence of experimental noise and the stochastic nature of quantum measurement.
Figure 4: Experimental Rydberg atoms array. a) Different Fourier-space order parameters (Eq. 15) highlighting the location of the checkerboard, star and striated phases across the phase space. The boundary ordered phase is located by means of the difference between the nearest-neighbor correlator for atoms at the edge and the bulk of the array. The insets showcase exemplary configurations of each phase on a smaller $9\times 9$ lattice, with white, purple and orange circles representing atoms in the |$\ket{g}$, |$\ket{r}$ and |$\ket{+_{x}}$ states, respectively. b) Latent representation learned by the cpVAE. For the two remaining active neurons, we present here the mean values (left) and variances (right) as output by the encoder, for input configurations across the phase space. We also show the absolute mean values (center), which showcase a finer resolution of the learned latent space.
@@
$H_{\mathrm{Rydberg}}=$ $\sum_{i<j}\frac{C_{6}}{||\mathbf{r_{i}}-\mathbf{r_{j}}||^{6}}\hat{n}_{i}\hat{n}_{j}-\Delta\sum_{i}\hat{n}_{i}$
@@
The phase space of the described system is illustrated in Fig. 4a. First, three different phases, namely the checkerboard, star and striated phases can be identified through the different Fourier-space order parameters [16]
~$\tilde{\mathcal{F}}(k_{1},k_{2})=\langle\mathcal{F}(k_{1},k_{2})+\mathcal{F}(k_{2},k_{1})\rangle/2$ (14)
@@
Each panel in Fig. 4a shows the particular parameter used to identify each phase (see panels’ titles), with insets showcasing exemplary configurations of each phase on a smaller $9\times 9$ lattice. Moreover, an edge-ordered phase can be also recognized by the difference between the nearest neigbour correlator of atoms at the edge and at the bulk of the array. For $R_{b}/a>1.5$ and between the edge-ordered and the star phase, an additional ordered phase was identified in Ref. [4].
@@
The left columns, corresponding to $\mu_{i}$, reveal that the latent dimensions are sufficient to discriminate between distinct quantum phases. For example, the regime in which the system energetically favors minimizing the number of atoms in the |$\ket{r}$ state ($\Delta/\Omega<0$) is characterized by low $\mu_{1}$ and high $\mu_{2}$ values. Then, the star phase shows low values of $\mu_{1}$, the striated phase low values in both components, and the checkerboard phase exhibits large $\mu_{1}$. By analyzing the absolute values $|\mu_{i}|$ (central columns of Fig. 4b), we found an even better resolved picture of the latent space, where the boundary-ordered phase is clearly highlighted for $|\mu_{1}|=0$.
The right columns of Fig. 4b display the variances $\sigma_{i}$ of the latent dimensions, which quantify the uncertainty associated with each neuron’s encoding. Lower values of $\sigma_{i}$ correspond to narrower posterior distributions, implying reduced variability in the sampled latent variable $z_{i}$ around its mean $\mu_{i}$. While the KL regularization term in the loss function encourages larger $\sigma_{i}$ to match the unit-variance prior $p(z)$, this pressure is counteracted by the reconstruction loss, which favours more precise sampling. In regions of the phase diagram where redundant or non-discriminative information is encoded, the model can tolerate increased variance without compromising reconstruction accuracy. This effect is visible, for example, at $\Delta/\Omega<0$, corresponding to nearly no atoms in the |$\ket{r}$ state, where $\sigma_{i}$ becomes large. By inspecting the variance across the latent dimensions, one gains insight into their relative importance in different regions of the latent space. In particular, the first latent shows reduced variance in the vicinity of the boundary-ordered and checkerboard phases, indicating its relevance in encoding these features. Conversely, the second latent exhibits small variance around the star and striated phases, suggesting its role in resolving these configurations. A similar study of the variances of the cpVAE trained on the spin model is provided in Appendix E, where we also further analyze their relationship to the informational content of the input configuration.
@@ IV Discussion
The proposed framework can also be integrated with complementary interpretable techniques. For instance, equipping the encoder with correlator neural networks—architectures designed to capture designated local and non-local correlations in the data [11]—would allow the latent variables to be analysed in tandem with explicit correlation features, yielding a richer, more transparent description of the underlying physics. Furthermore, as demonstrated in Section III.1, symbolic-regression methods offer a promising approach to extracting compact analytical expressions from the learned latents. The dimensionality and structure revealed by the cpVAE can in turn serve as informed priors for such regression algorithms, accelerating their convergence and enhancing the interpretability of the resulting formulae.
@@ V Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [SFB BeyondC F7102, 10.55776/F71]. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This work was also supported by the European Union (ERC Advanced Grant, QuantAI, No. 101055129). The views and opinions expressed in this article are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council - neither the European Union nor the granting authority can be held responsible for them. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ERCAP0032002.
@@ VI Code availability
The code needed to reproduce the results presented in this work is available in Ref. [29]. The datasets for the NNN-TFIM and the LR-TFIM were created by means of NetKet [38, 39], which is based on Jax [40] and MPI4Jax [41]. Netket has also been used for implementing the dense ARNN used as decoder of the cpVAE on the spin models. The implementation of the TC loss (Eq. 7) was done based on the implementation in [42]. The neural networks have been implemented using Flax [43].
@@ Appendix A Neural network architectures
A.1 Encoder
For the VAEs trained on both spin models in  Section III.1, a convolutional neural network was used as the encoder. The details of its architecture are presented in Table 1. The same architecture is used for both dVAE and cpVAE.
@@ A.1 Encoder
        Dense $B\times 5$
Table 1: Convolutional encoder. Details on the encoder architecture used for the NNN-TFIM and the LR-TFIM. $B$ is the batch size, $N$ the size of the spin configurations and $k_{s}$ the kernel size. The latent space dimension was chosen to be 5. The same architecture is used for the dVAE and the cpVAE.
On the other hand, for the cpVAE trained on snapshots of Rydberg atom arrays (Section III.2), we used a transformer architecture for its encoder [44]. Further details are presented in Table 2.
@@
Table 2: Transformer encoder. Details on the encoder architecture used for the cpVAE trained on snapshots of Rydberg atoms. $B$ in the batch size, $N$ is the number of atoms. A dimension of $d_{\mathrm{model}}=8$ was used. The latent space dimension was chosen to be 5.
@@ A.2 Autoregressive decoder
As mentioned in the main text (see Section II), autoregressive neural network are used to model the conditional probabilities of each spin values
@@
A common approach to render a neural network autoregressive is to impose a masking scheme that prevents access to future elements of the input sequence. For models trained on spin models (Section III.1), we use as the decoder of the cpVAE a standard feedforward neural network (FFNN), with the masking being achieved by simply imposing triangular weights. For the numerical simulations, we use 3 hidden layers with 80 hidden neurons and selu activation functions [45]. This design choice allowed us to employ the same neural network architecture for the decoder of the dVAE, omitting only the autoregressive masking used in the cpVAE. This ensures that comparisons between the two models remain as fair as possible.
For the Rydberg atom arrays, as inspired by Ref. [46], we use the typical transformer architecture [44] for the decoder of the cpVAE, as detailed in Table 3. The latent representation output by the encoder is fed into the transformer as context, following the typical context-based attention mechanism.
@@
Table 3: Transformer decoder. Details on the decoder architecture used for cpVAE trained on snapshots of the Rydberg atoms. $B$ is the batch size and $N$ the number of atoms. For the model presented in the main text, we fixed $d_{\mathrm{model}}=8$ and the latent space dimension to $d_{\mathrm{latent}}=5$. Not shown here, the context vector used in the context-based attention are the latent variables $z_{i}$, after a linear layer expending their dimension to $B\times d_{\mathrm{latent}}\times d_{\mathrm{model}}$
@@ B.1 Training hyperparameters
For the experimental system of Rydberg atoms, we had access to $250$ snapshots taken for each set of parameters $(R_{b}/a,\Delta/\Omega)$ In particular, the parameters were
@@
We then optimized the neural network with the adabelief optimizer [47], with default parameters and a leaning rate of 0.001.
B.2 Training curves
@@
For the dVAE, the reconstruction loss corresponds to the mean squared error (MSE). Across both spin models, this loss decreases rapidly in the early training stages and subsequently plateaus. This typically indicates that the model has insufficient capacity or expressivity to fully learn the data distribution [48]. In our case, as shown in the main text, this model was not able to approximate the quantum probability distribution and capture the correlations. As discussed in the main text, in order to ensure a fair comparison between the dVAE and cpVAE, the dVAE was trained without KL regularization. Additionally, its latent dimension was fixed to match the expected number of relevant latent factors, as found by the cpVAE and previously known for the presented benchmark cases. Under these conditions, we observe that the variances $\log\sigma_{i}$ decrease rapidly for both spin models (see more about this below).
@@
Figure 6: Training: LR-TFIM. Training of a) the dVAE and b) the cpVAE seen through the reconstruction loss $\mathcal{L}_{\mathrm{reconstr.}}$ and the $\log$ of the variances $\sigma_{i}$ outputt by the encoder. For the dVAE the reconstruction loss is the mean square error (MSE) and for the cpVAE it is the first term in Eq. 7.
@@
Figure 7: Training: Rydberg atoms. Training of the cpVAE seen through the reconstruction loss $\mathcal{L}_{\mathrm{reconstr.}}$ (left) and the $\log$ of the variances $\sigma_{i}$ outputt by the encoder (right).
@@ Appendix C NNN-TFIM: Generalization abilities
The generalization capabilities of the dVAE and cpVAE were also investigated on the NNN-TFIM using a restricted training set. Specifically, both models were trained without any spin configuration coming from the region $0.750.4<h<$ 0.75 of the parameter space. We then evaluated the generation performance of both models across the entire phase space, including the excluded region. The reconstructed values of the nearest-neighbor correlator are presented in Fig. 8a) and b) for the dVAE and cpVAE, respectively. Notably, the quality of the reconstructed correlators in the previously unseen region remains consistent with that of the training regions. This indicates that both architectures exhibit robust generalization to regions of the phase space that were not included in the training set.
Figure 8: Generalization abilities: NNN-TFIM. Next-nearest-neigbour correlator of the spin configuration generated by a) the dVAE and b) the cpVAE. The training of both VAE were performed without spin configurations coming from the ground state of Hamiltonians having $0.750.4<h<$ 0.75, region inside the orange rectangle. Then they are tested on the entire space.
@@ Appendix D LR-TFIM: Sampling from the latent space
Figure 9: Generation: LR-TFIM. To further interpret the structure learned by the cpVAE, a scan is performed over the latent variables $z_{1,2}$. Then, for each values of the latter, samples are generated using the decoder of the cpVAE. Left is the magnetization of these samples, middle is the $\beta$ exponent and right is the correlatiors $\langle Z^{i}Z^{i+1}\rangle$.
To further interpret the structure learned by the cpVAE, we leverage its generative nature to probe the contents of the latent space. A linear sweep is performed over each latent dimension independently, and for every point along the scan, 1000 spin configurations are sampled from the generative decoder. For each set of generated configurations, we compute the magnetization, the correlation exponent $\beta$ and the nearest neigbour correlator $\langle Z^{i}Z^{i+1}\rangle$. The results are presented in Fig. 9. This generative probing again highlights that the two latent dimensions encode distinct physical features: the first latent neuron predominantly controls the magnetization of the output configurations, while the second mainly governs their correlation. Notably, for certain latent values, the model generates spin configurations with an apparent $\beta<0$. These correspond to atypical states that exhibit very weak or inverted correlation patterns, in which nearest-neighbour correlations vanish and increase slightly with distance. In these cases, the power-law fit used to estimate $\beta$ becomes unreliable.
Appendix E Study of the latent neuron variances for spin models benchmarks

### v3

@@ III.1.2 LR-TFIM
As a second benchmark, we consider the long-range transverse field Ising model (LR-TFIM), which has been recently performed experimentally via trapped-ion simulators [hauke2013spread, jurcevic2014quasiparticle, li2023probing]. This model is a modification of the conventional TFIM in which spin–spin interactions decay algebraically with distance. Its Hamiltonian is given by
@@
In contrast, the latent space learned by the cpVAE (Fig. 3c) exhibits a markedly richer structure. The first latent neuron, $\mu_{1}$, correlates strongly with a magnetization-like observable (see Appendix D for more details), which is consistent with it capturing the dominant factor of variation in the data. The second, $\mu_{2}$, encodes a distinct feature that highlights the presence of an intermediate phase emerging at low values of $\alpha$ and $h$. To investigate this further, we applied symbolic regression [crammer2023interpretable] to the latent values of $\mu_{2}$, seeking a compact analytical expression $f(x)$ that best describes its behavior given the corresponding spin configuration $x$ (see Appendix E for more details). This approach suggests that the neuron was encoding a quantity similar to the structure factor [baez2020dynamical]
@@
As shown in Fig. 3d, the structure factor, evaluated here at fixed momentum $k=0.8$ and spatial reference site $i=0$, highlights the same region of the phase diagram as captured by the latent variable $\mu_{2}$. From a physical perspective, the structure factor reveals a distinct intermediate regime between the ferromagnetic phase and the high $\beta$ paramagnetic regime. This feature, visible only at the chain boundaries (i.e., $i=0$ and $i=N$), suggests the emergence of a boundary-induced phase, likely arising from finite-size and open-boundary effects [vodola2015long]. That the cpVAE identifies this behaviour without supervision underscores its potential as a tool for uncovering nontrivial structure in many-body systems and guiding the formulation of new hypotheses.
@@ III.2 Experimental Rydberg Atoms
[…]
@@
[…]
@@
The phase space of the described system is illustrated in Fig. 4a. First, three different phases, namely the checkerboard, star and striated phases can be identified through the different Fourier-space order parameters [ebadi2021quantum]
$\tilde{\mathcal{F}}(k_{1},k_{2})=\langle\mathcal{F}(k_{1},k_{2})+\mathcal{F}(k_{2},k_{1})\rangle/2$ (14)
@@
Each panel in Fig. 4a shows the particular parameter used to identify each phase (see panels’ titles), with insets showcasing exemplary configurations of each phase on a smaller $9\times 9$ lattice. Moreover, an edge-ordered phase can be also recognized by the difference between the nearest neighbor correlator of atoms at the edge and at the bulk of the array. For $R_{b}/a>1.5$ and between the edge-ordered and the star phase, an additional ordered phase was identified in Ref. [miles2023machine].
@@
[…]
@@ IV Discussion
The proposed framework can also be integrated with complementary interpretable techniques. For instance, equipping the encoder with correlator neural networks—architectures designed to capture designated local and non-local correlations in the data [miles2021correlator]—would allow the latent variables to be analysed in tandem with explicit correlation features, yielding a richer, more transparent description of the underlying physics. Furthermore, as demonstrated in Section III.1, symbolic-regression methods offer a promising approach to extracting compact analytical expressions from the learned latents. The dimensionality and structure revealed by the cpVAE can in turn serve as informed priors for such regression algorithms, accelerating their convergence and enhancing the interpretability of the resulting formulae.
@@ V Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [SFB BeyondC F7102, DOI: 10.55776/F71; WIT9503323, DOI: 10.55776/WIT9503323]. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This work was also supported by the European Union (ERC Advanced Grant, QuantAI, No. 101055129). The views and opinions expressed in this article are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council - neither the European Union nor the granting authority can be held responsible for them.
@@ VI Code availability
The code needed to reproduce the results presented in this work is available in Ref. [github]. The datasets for the NNN-TFIM and the LR-TFIM were created by means of NetKet [filippo2022netket, filippo2022codebase], which is based on Jax [james2018jax] and MPI4Jax [dion2021mpi4jax]. Netket has also been used for implementing the dense ARNN used as decoder of the cpVAE on the spin models. The implementation of the TC loss (Eq. 7) was done based on the implementation in [dubois2019disentangling]. The neural networks have been implemented using Flax [heek2024flax].
@@ Appendix A Neural network architectures
A.1 Encoder and latent space
For the VAEs trained on both spin models in  Section III.1, a convolutional neural network was used as the encoder. The details of its architecture are presented in Table 1. The same architecture is used for both dVAE and cpVAE. On the other hand, for the cpVAE trained on snapshots of Rydberg atom arrays (Section III.2), we used a transformer architecture for its encoder [vaswani2017attention]. Further details are presented in Table 2.
In terms of the latent space, we consider for all cpVAE a latent dimension of $d_{z}=5$. We remind the reader that through training and minimization of Eq. 1, the model will autonomously noise out unnecessary neurons until it finds a minimal representation (i.e., 1 for the NNN-TFIM, 2 for LR-TFIM and 2 for the Rydberg atom dataset). We hence recommend training VAEs with large latent dimensions $d_{z}$, typically much larger than the expected degrees of freedom underlying the data. The model will then autonomously determine the optimal dimension.
@@ A.1 Encoder and latent space
      Dense $B\times(d_{z}=5)$
Table 1: Convolutional encoder. Details on the encoder architecture used for the NNN-TFIM and the LR-TFIM. $B$ is the batch size, NN the size of the spin configurations and $k_{s}$ the kernel size. The latent space dimension was chosen to be 5. The same architecture is used for the dVAE and the cpVAE.
@@
Table 2: Transformer encoder. Details on the encoder architecture used for the cpVAE trained on snapshots of Rydberg atoms. $B$ in the batch size, NN is the number of atoms. A dimension of $d_{\mathrm{model}}=8$ was used. The latent space dimension was chosen to be 5.
@@ A.2 Autoregressive decoder
As mentioned in the main text (see Section II), autoregressive neural networks are used to model the conditional probabilities of each spin values
@@
A common approach to render a neural network autoregressive is to impose a masking scheme that prevents access to future elements of the input sequence. For models trained on spin models (Section III.1), we use as the decoder of the cpVAE a standard feedforward neural network (FFNN), with the masking being achieved by simply imposing triangular weights. For the numerical simulations, we use 3 hidden layers with 80 hidden neurons and selu activation functions [klambauer2017self]. This design choice allowed us to employ the same neural network architecture for the decoder of the dVAE, omitting only the autoregressive masking used in the cpVAE. This ensures that comparisons between the two models remain as fair as possible.
Due to the higher complexity of the Rydberg atom datasets, we consider here a more powerful decoder based on the transformer architecture [vaswani2017attention] (see details in Table 3). The latent representation output by the encoder is fed into the transformer as context, following the typical context-based attention mechanism. This autoregressive approach requires an ordering of the 2D lattice atoms to be specified. We chose to use a snake ordering, which is commonly used for 2D systems [lange2024from, fitzek2024rydberggpt].
@@
Table 3: Transformer decoder. Details on the decoder architecture used for cpVAE trained on snapshots of the Rydberg atoms. $B$ is the batch size and $N$ the number of atoms. For the model presented in the main text, we fixed $d_{\mathrm{model}}=8$ and the latent space dimension to $d_{\mathrm{latent}}=5$. Not shown here, the context vector used in the context-based attention are the latent variables $z_{i}$, after a linear layer expanding their dimension to $B\times d_{\mathrm{latent}}\times d_{\mathrm{model}}$
@@ B.1 Hyperparameters
For the experimental system of Rydberg atoms, we had access to $250$ snapshots taken for each set of parameters $(R_{b}/a,\Delta/\Omega)$ . In particular, the parameters were
@@
We then optimized the neural network with the adabelief optimizer [zhuang2020adabelief], with default parameters and a learning rate of 0.001. The training stops when the reconstruction loss and the variance of the latent variables converge.
B.2 Training
@@
For the dVAE, the reconstruction loss corresponds to the mean squared error (MSE). Across both spin models, this loss decreases rapidly in the early training stages and subsequently plateaus. This typically indicates that the model has insufficient capacity or expressivity to fully learn the data distribution [smith2018disciplinedapproachneuralnetwork]. In our case, as shown in the main text, this model was not able to approximate the quantum probability distribution and capture the correlations.
In order to ensure a fair comparison between the dVAE and cpVAE, the dVAE was trained without KL regularization. Its latent dimension was fixed to match the expected number of relevant latent factors, as found by the cpVAE and previously known for the presented benchmark cases. Under these conditions, we observe that the variances $\log\sigma_{i}$ decrease rapidly for both spin models.
When training without KL regularization and restricting the latent dimension, our goal is to design the fairest possible comparison between the dVAE and the proposed cpVAE. As discussed in the main text, the main purpose of the KL divergence term is to minimize the number of active neurons in the latent space to create a minimal representation of the input data. This is the main challenge in VAE training and typically depends on good hyperparameterisation of the weights in Eq. 7.
With the proposed approach, this objective (learning a minimal latent representation) is achieved directly by constraining the latent dimensionality of the dVAE. Introducing the KL term would impose additional constraints on the latent space of the dVAE and could reduce the expressiveness of the latent representation. Furthermore, tuning the KL weight introduces additional hyperparameters, which complicates the comparison and risks introducing bias in favor of one model. By avoiding the KL term, we consider the ’best possible’ scenario for the dVAE, which still fails to match the representation power of the cpVAE.
Meanwhile, the KL term, and more importantly, its TC decomposition (see Eq. 6), helps the model find more interpretable representations by promoting uncorrelated latent neurons, a feature commonly known as disentanglement. However, it should be noted that the aim here is not to compare the structure of the latent representations, but rather to demonstrate that the dVAE cannot learn parts of the phase space due to its deterministic nature, in contrast to the cpVAE.
@@
Figure 6: Training: LR-TFIM. Training of a) the dVAE and b) the cpVAE seen through the reconstruction loss $\mathcal{L}_{\mathrm{reconstr.}}$ and the $\log$ of the variances $\sigma_{i}$ output by the encoder. For the dVAE the reconstruction loss is the mean square error (MSE) and for the cpVAE it is the first term in Eq. 7.
@@
Figure 7: Training: Rydberg atoms. Training of the cpVAE seen through the reconstruction loss $\mathcal{L}_{\mathrm{reconstr.}}$ (left) and the $\log$ of the variances $\sigma_{i}$ output by the encoder (right).
@@ Appendix C NNN-TFIM: Generalization abilities
The generalization capabilities of the dVAE and cpVAE were also investigated on the NNN-TFIM using a restricted training set. Specifically, both models were trained without any spin configuration coming from the region $0.4<h<0.75$ of the parameter space. We then evaluated the generation performance of both models across the entire phase space, including the excluded region. The reconstructed values of the nearest-neighbor correlator are presented in Fig. 8a and b for the dVAE and cpVAE, respectively. Notably, the quality of the reconstructed correlators in the previously unseen region remains consistent with that of the training regions. This indicates that both architectures exhibit robust generalization to regions of the phase space that were not included in the training set.
Figure 8: Generalization abilities: NNN-TFIM. Next-nearest-neighbor correlator of the spin configuration generated by a) the dVAE and b) the cpVAE. The training of both VAE were performed without spin configurations coming from the ground state of Hamiltonians having $0.4<h<0.75$, region inside the orange rectangle. Then they are tested on the entire space.
@@ Appendix D LR-TFIM: Sampling from the latent space
Figure 9: Generation: LR-TFIM. To further interpret the structure learned by the cpVAE, a scan is performed over the latent variables $z_{1,2}$. Then, for each values of the latter, samples are generated using the decoder of the cpVAE. Left is the magnetization of these samples, middle is the $\beta$ exponent and right is the correlators $\langle Z^{i}Z^{i+1}\rangle$.
To further interpret the structure learned by the cpVAE, we leverage its generative nature to probe the contents of the latent space. A linear sweep is performed over each latent dimension independently, and for every point along the scan, 1000 spin configurations are sampled from the generative decoder. For each set of generated configurations, we compute the magnetization, the correlation exponent $\beta$ and the nearest neighbor correlator $\langle Z^{i}Z^{i+1}\rangle$. The results are presented in Fig. 9. This generative probing again highlights that the two latent dimensions encode distinct physical features: the first latent neuron predominantly controls the magnetization of the output configurations, while the second mainly governs their correlation. Notably, for certain latent values, the model generates spin configurations with an apparent $\beta<0$. These correspond to atypical states that exhibit very weak or inverted correlation patterns, in which nearest-neighbor correlations vanish and increase slightly with distance. In these cases, the power-law fit used to estimate $\beta$ becomes unreliable.
Appendix E Symbolic regression methods to extract analytical expressions
As mentioned in the main text, for the LR-TFIM, the second active latent neuron of the cpVAE highlighted the presence of an intermediate phase at low values of $\alpha$ and $h$. To further investigate the physical nature of this emergent region, we applied symbolic methods. In particular, we used symbolic regression (SR), implemented via the open-source PySR package [crammer2023interpretable], to extract closed-form analytical relations that describe this additional phase in the latent representation.
PySR is based on a multi-population evolutionary algorithm that evolves sets of candidate expressions through repeated mutation, crossover and simplification. The algorithm refines these expressions iteratively using a loop comprising the following three steps: (i) new candidates are evolved via symbolic mutations; (ii) expressions are simplified using algebraic rules to avoid redundant complexity; and (iii) scalar constants are optimized using gradient-based solvers. This loop enables PySR to discover compact analytical relations containing unknown real parameters, which is crucial for scientific applications.
We design the SR to find a symbolic expression $f(\mathbf{x})$ that distinguishes the spin configurations $\mathbf{x}$ that come from the additional phase (inducing $\mu_{2}>0.35$) from those outside. The best expressions we recovered are dominated by two-body correlator terms with alternating signs. For example, one extracted expression (truncated) was
$f(\mathbf{x})=-x_{0}x_{1}-2x_{0}x_{1}-x_{0}x_{12}-x_{0}x_{14}-x_{0}x_{16}-4x_{0}x_{17}+2x_{0}x_{12}+x_{0}x_{19}+4.5x_{0}x_{2}-3.2x_{0}x_{3}+x_{0}x_{8}+$...
Although studying these expressions led us to the correct structure factor parameter, these fits were neither robust nor easily interpretable, hence we have omitted them from the main results (see Eq. 11). Nevertheless, as mentioned in the conclusion, we regard the combination of learned latent representations with symbolic regression as a promising direction, although we leave its proper integration within the VAE pipeline for future work.
Appendix F Study of the latent neuron variances for spin models benchmarks